# All-fiber microendoscopic polarization sensing at single-photon level aided by deep-learning

### v2

@@ Abstract
The polarization of light conveys crucial information about the spatial ordering and optical properties of a specimen. However, precise polarization measurement in challenging conditions, including constrained spaces, low light levels, and high-speed scenarios, remains a severe challenge. Addressing this problem, we introduce a real-time polarization measurement method that is accurate down to a single-photon level and provides complete information about the polarization state. Free of moving components, the polarization sensor utilizes a few-mode fiber followed by a fiber array and a detector array. The calibration of the sensor relies on a neural network yielding unprecedented accuracy across all polarization states, including partially polarized light. We validate the approach by visualizing the polarization structure of biological specimens and the birefringent USAF test. Our method offers an efficient and reliable solution for real-time polarization sensing and microendoscopy under low-light conditions.
journal: opticajournal
@@ 1 Introduction
From ellipsometry to optical communications, from microscopy to sensing, light polarization represents a fundamental element across various scientific disciplines. In particular, the polarization of light emitted by an optically anisotropic specimen carries essential information about its material structure and optical properties. Consequently, conducting a thorough analysis of such specimens demands precise characterization of the emitted light polarization. While relatively straightforward in a controlled environment, accurate polarization measurement under challenging conditions remains a significant obstacle.
Low light intensities, representing the first challenging condition, come with a reduced signal-to-noise ratio, which complicates extracting meaningful polarization information from data burdened with environmental and shot noise. Similarly, conducting high-speed measurements, as the second condition, is accompanied by low signal levels within the acquisition time. While increasing the incident light intensity can help, it is limited by our ability to deliver the optical power to the particular sensing area and, most ultimately, by the photodamage threshold of the sample. Incorporating single-photon detectors can substantially improve the accuracy of protocols relying on the number of detected photons for polarization state reconstruction. However, another drawback arises as these detectors are relatively large and cannot be efficiently integrated. This issue becomes particularly problematic when faced with the third condition - constrained space. Such spaces not only impose restrictions on the overall size of the measurement setup but also limit the placement and alignment of optical components. Therefore, these challenging conditions pose a severe setback, especially for measurements during invasive medical procedures, such as microendoscopy, or internal inspections of materials and products. We address these problems by developing a polarization sensor based on light propagation in disordered media.
@@
Coherent light propagating within a disordered medium is subjected to multimode interference, generating a granular speckle pattern. Despite the apparent degrading effect, the pattern retains encoded information about the incident light [1, 2, 3, 4]. Analyzing the light propagation in disordered media and employing suitable post-processing techniques, one can study the disordering effects of the media [5, 6], estimate and classify input images [7, 8, 9, 10, 11, 12], learn to modulate the input for projecting a targeted image on the output [13], reconstruct shapes of propagated ultrashort laser pulses [14, 15], compute and communicate [16, 17, 18, 19, 20], and even make an accurate spectrometer or wavemeter [21, 22]. Recently, the speckle pattern has proved to preserve a great deal of information about the polarization state of the incident light [23, 24]. Although additional studies successfully advancing the application of disordered-media propagation have been conducted [25, 26, 27, 28], developing a compact polarimeter capable of rapid measurements with single-photon sensitivity continues to be an open problem.
2 Experimental and data processing setup
The speckle-based polarization sensor depicted in Fig. 1 employs a common few-mode fiber to characterize the input polarization state. The sensor collects light with a front face of a 5 cm long step-index SMF28 fiber with a numerical aperture of 0.14 and a core diameter of 8.2 $\mu$m. While the motivation behind choosing this specific fiber type is to minimize the sensing area and keep the speckle structures large, we emphasize that this preference is optional. Virtually any few-mode or multimode fiber, including significantly shorter variants, can be used as a substitute. The main requirement for the fiber is to support several interfering modes necessary for creating a polarization-dependent speckle pattern at the output. Following a 5mm free-space propagation, we collect discrete samples of this interference pattern using a fiber array comprising seven gradient multimode fibers, each with a 62.5 $\mu$m core diameter, forming a 375 $\mu$m diameter honeycomb pattern. Each fiber then propagates the collected intensity signal to an independent single-photon avalanche diode, recording the number of detection events. After normalizing the measured counts, this setup associates the sampled polarization state with the corresponding relative frequencies of each detector, which we term a count distribution. For more experimental details, see Sec. S1 in Supplement 1.
@@ 2 Experimental and data processing setup
We used the dataset to train a fully-connected deep neural network for reconstructing the polarization state given the associated count distribution, see Sec. S1 for details. The optimized and fully-trained network was evaluated on a test set of previously unobserved polarization states, also acquired from the experimental setup. We achieved the unprecedented average infidelity of $8\times 10^{-4}$ with a $[5\times 10^{-5},2\times 10^{-3}]$ confidence interval using the 10th and 90th percentile. These results reflect the highest achievable measurement accuracy as they approach the fidelity of our polarization state preparation. Furthermore, we conducted additional validation of our results using a 1951 USAF test, as detailed in Sec. S2 of Supplement 1. This supplemental analysis provides further insight into the overall performance of our approach.
3 Performance evaluation
@@
Furthermore, let us discuss the long-term stability of our approach and its resistance to environmental effects. Despite using a 5cm long few-mode fiber in an open environment without elaborated fixation, the polarization sensor shows sufficient stability for a week without additional calibration or network retraining. For in-situ and microendoscopic applications, the fiber is typically inserted in a cannula, preventing bending and micro deformations. The stability can be further enhanced using a few-mm or sub-mm long fiber encased within a ceramic or glass ferrule tip. The millimeter-long fiber provides sufficient intermodal coupling, and the encasing fixes the fiber and its position relative to the fiber array. With these modifications, we expect an improvement in stability by up to two orders of magnitude. In summary, our polarization sensor represents an outstanding instrument for efficient and accurate real-time and high-speed polarimetry applications.
We first demonstrate the sensing ability of our approach by conducting a polarization-sensitive scan of dense connective tissue (AmScope PS25W). While the primary application of our sensor significantly differs from polarization imaging, this optically anisotropic specimen is famously known in the polarization microscopy field. Visualizing its typical structures further underscores the accuracy of the presented polarimetric sensor. Using the experimental arrangement outlined in Sec. 2 and Supplement 1 Sec. S1, we replaced the calibration polarization preparation stage with the tissue specimen. The few-mode fiber positioned proximal to the tissue scanned over its area, capturing the induced polarization transformation. The reconstructed polarization states were expressed as three Bloch parameters, i.e., renormalized Stokes parameters [3], and visualized as a false-colored RGB image. In Fig. 3, we present a side-by-side comparison of this polarization scan with an intensity image captured from an identical tissue region. This intensity image was obtained using a monochrome camera (ImagingSource DMK 23U274) with a pixel size of $4.4\times 4.4\leavevmode\nobreak\ \mu$m. We illuminated the specimen with 0.8 $\mu$m light and projected the resulting image onto the camera through a 25.4 mm focal-length lens, providing an estimated fivefold magnification. As illustrated in Fig. 3, the polarization sensor reveals underlying polarization structures that remain concealed in the intensity profile. Refer to Fig. S3 of Supplement 1 for a comparison with the corresponding image acquired using a stand-alone polarization microscope.
@@
Figure 3: Visualization of dense connective tissue: (left) intensity image and (right) a scan using the all-fiber polarization sensor. The resulting Bloch parameters are represented as an RGB false-colored image. The three highlighted pixels characterize the purple and green segments in the polarization structure alongside a reference background polarization. Their respective Bloch parameters are $\text{A}=(0.34,-0.43,0.82),$ $\text{B}=(-0.90,0.05,0.06),$ and $\text{C}=(-0.04,-0.18,0.98).$
@@
As the dense connective tissue is a static specimen, the previous polarization scan can be acquired using advanced imaging polarimeters based on rotating waveplates and a CCD/CMOS camera. However, a polarization measurement using wave plates is an insufficient method for characterizing a rapidly changing polarization state induced by moving or evolving specimens. These scenarios require significantly faster approaches that can be performed in a single-shot regime. We demonstrate the feasibility of our sensor for such applications by conducting a polarization measurement to characterize the floating diatom. These microscopic unicellular organisms comprise a complex inner structure that can exhibit anisotropic properties. By continuously measuring the polarization state, we characterized the polarization transition caused by a moving diatom within the same experimental arrangement as in the previous application. The evolution of the Bloch parameters corresponding to this dynamic process is depicted in Fig. 4 with red lines. Compared to the blue lines of background polarization, the diatom properties are visible in the modulation of Bloch parameters. One could count the number of living specimens, estimate their size and shape, or even classify the detected organisms based on their birefringent properties. The sensor can also find applications for in-situ material inspection, local strain analysis, and micro crystal growth monitoring. These showcases underscore the versatility of our polarization sensor for various sensing applications and its valuable contributions to low-light high-speed polarimetry.
@@ 4 Conclusion
We have developed a single-shot polarization sensor based on light propagation through a short piece of a few-mode fiber, resulting in a polarization-dependent speckle pattern. The defining aspect of our approach is the sparse sampling of this pattern using a fiber array instead of capturing the whole intensity profile. This characteristic allows the use of any photonic detectors, including high-speed or single-photon detectors, enabling fast response and operation at low light intensities. The implementation of deep learning algorithms in the system ensures an accurate reconstruction, providing complete information about the polarization state, including partial polarization, with an infidelity of $8\times 10^{-4}$. Complemented by the high operation speed, it allows for both real-time and high-repetition-rate polarization sensing with remarkable precision. The all-fiber sensor, featuring a compact design with no moving components, is particularly well-suited for reducing the invasiveness of biomedical procedures in constrained spaces and under low-light conditions. Furthermore, we can enhance the spatial resolution of our sensor by tapering the tip of the few-mode fiber. This method could even provide comparable resolution as near-field scanning optical microscopy while preserving the single-photon polarization sensitivity. In the view of optically-assisted machine learning, the random propagation in multimode fiber followed by a computational neural network represents an instance of an extreme learning machine or reservoir computing [19, 32]. By demonstrating the single-photon sensing capability of these architectures, we have provided the foundation for decreasing energy consumption in optical sensing and computing. In summary, our polarization sensor significantly advances the field, offering a combination of compactness, accuracy, low light sensitivity, and high operational speed that opens up new possibilities for highly accurate polarization sensing across various applications.
[…]
Funding This work was supported by the Czech Science Foundation (project 21-18545S), Ministry of Education, Youth, and Sports of the Czech Republic (project OP JAC CZ.02.01.01/00/23_021/0008790). MB and DV acknowledge the support by Palacký University (projects IGA-PrF-2023-006 and IGA-PrF-2024-008).
[…]
Acknowledgments We acknowledge the use of cluster computing resources provided by the Department of Optics, Palacký University Olomouc. We thank Jan Provazník for maintaining the cluster and providing support. We also thank Josef Hloušek for developing a 3D model shown in Fig. 1 and Jaromír Běhal for a fruitful discussion on birefringent biological samples.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data Availability Statement The code and data that support the findings of this study are publicly available on GitHub [33].
Supplemental Document
1 Detailed sensor description
@@ 1.1 Experimental setup
In addition to the information provided in the main text, we include a detailed scheme of our all-fiber polarization sensor, illustrated in Fig. S1. This scheme comprises three configurations. In the first configuration, the twisted nematic liquid crystal device [1] induces polarization transformation on the attenuated coherent signal generated by a fiber-coupled continuous 810 nm semiconductor laser diode (QPhotonics QFLD). We independently verified the accurate operation of the liquid crystal device in the second configuration by replacing it with a reference polarimeter based on rotating wave plates. The third configuration represents the sensor application and comprises the birefringent specimen under investigation. In all cases, the modified signal is collected by the 5 cm long SMF28 few-mode fiber. The number of modes supported by this fiber is characterized in Subsection 1.2.
@@
Figure S1: A detailed visualization of the experimental setup for the all-fiber sensor. The twisted nematic liquid crystal device, labeled as (a), is present only during dataset acquisition for deep neural network training. In the verification stage, depicted in (b), a reference polarimeter based on wave plates is employed to validate the accurate operation of the liquid crystal device. Finally, this segment is replaced with a polarization-modulating specimen (c) during the USAF test, dense connective tissue, and diatom measurements.
@@ 1.2 Supported fiber modes analysis
$V=\frac{2\pi}{\lambda}r_{\text{core}}\text{NA},$ (S1)
where $\lambda=810\leavevmode\nobreak\ \text{nm}$ is the wavelength of propagated light, $r_{\text{core}}=4.1\leavevmode\nobreak\ \mu\text{m}$ is the fiber core radius, and $\text{NA}=0.14$ is the numerical aperture. The calculated value of $V=4.45>2.405$ indicates that the fiber supports more than one mode per polarization direction. Subsequently, the number of modes $M$ is approximately given by the relation
$M\approx\frac{V^{2}}{2}=9.9,$ (S2)

### v3

@@ Abstract
The polarization of light conveys crucial information about the spatial ordering and optical properties of a specimen. However, precise polarization measurement in challenging conditions, including constrained spaces, low light levels, and high-speed scenarios, remains a severe challenge. Addressing this problem, we introduce a real-time polarization measurement method that is accurate down to a single-photon level and provides complete information about the polarization state. Free of moving components, the polarization sensor utilizes a short rigid piece of few-mode fiber followed by a fiber array and a detector array. The calibration of the sensor relies on a neural network yielding unprecedented accuracy across all polarization states, including partially polarized light. We validate the approach by visualizing the polarization structure of biological specimens and the liquid crystal polymer sample (birefringent USAF test). Our method offers an efficient and reliable solution for real-time polarization sensing and microendoscopy under low-light conditions.
keywords:
light polarization, all-fiber sensor, in-situ and microendoscopic measurement, single-photon and low-light sensing, deep learning
@@ 1 Introduction
From imaging [1, 2] to sensing [3], from ellipsometry [4, 5] to optical communications [6, 7], light polarization represents a fundamental element across various scientific disciplines. In particular, the polarization of light emitted by an optically anisotropic specimen carries essential information about its material structure and optical properties. Consequently, conducting a thorough analysis of such specimens demands precise characterization of the emitted light polarization. While relatively straightforward in a controlled environment, accurate polarization measurement under challenging conditions remains a significant obstacle.
Low light intensities, representing the first challenging condition, come with a reduced signal-to-noise ratio, which complicates extracting meaningful polarization information from data burdened with environmental and shot noise. Similarly, conducting high-speed measurements, as the second condition, is accompanied by low signal levels within the acquisition time. While increasing the incident light intensity can help, it is limited by our ability to deliver the optical power to the particular sensing area and, most ultimately, by the photodamage threshold of the sample. Incorporating single-photon detectors can substantially improve the accuracy of protocols relying on the number of detected photons for polarization state reconstruction. However, another drawback arises as these detectors are relatively large and cannot be efficiently integrated. This issue becomes particularly problematic when faced with the third condition - constrained space. Such spaces not only impose restrictions on the overall size of the measurement setup but also limit the placement and alignment of optical components. Therefore, these challenging conditions pose a severe setback, especially for measurements during invasive medical procedures, such as microendoscopy [8, 9], environmental monitoring [10, 11, 12], or internal/in-situ inspections of materials [13, 14]. We address these problems by developing a polarization sensor based on light propagation in disordered media.
Coherent light propagating within a disordered medium is subjected to multimode interference, generating a granular speckle pattern. Despite the apparent degrading effect, the pattern retains encoded information about the incident light [15, 16, 17, 18, 19]. Analyzing the light propagation in disordered media and employing suitable post-processing techniques, one can study the disordering effects of the media [20, 21], estimate and classify input images [22, 23, 24, 25, 26, 27, 9], learn to modulate the input for projecting a targeted image on the output [28], reconstruct shapes of propagated ultrashort laser pulses [29, 30], compute and communicate [31, 32, 33, 34, 35], and even make an accurate spectrometer or wavemeter [36, 37]. Recently, the speckle pattern has proved to preserve a great deal of information about the polarization state of the incident light [38, 39]. Although additional studies successfully advancing the application of disordered-media propagation have been conducted [40, 41, 42, 43], developing a compact polarimeter capable of rapid measurements with single-photon sensitivity continues to be an open problem.
@@
2 Results and Discussion
2.1 Experimental and data processing setup
The speckle-based polarization sensor depicted in Fig. 1 employs a common few-mode fiber to characterize the input polarization state. The sensor collects light with a front face of a 5 cm long step-index SMF28 fiber with a numerical aperture of 0.14 and a core diameter of 8.2 $\mu$m. While the motivation behind choosing this specific fiber type is to minimize the sensing area and keep the speckle structures large, we emphasize that this preference is optional. Virtually any few-mode or multimode fiber, including significantly shorter variants, can be used as a substitute. The main requirement for the fiber is to support several interfering modes necessary for creating a polarization-dependent speckle pattern at the output. Following a 5mm free-space propagation, we collect discrete samples of this interference pattern using a fiber array comprising seven gradient multimode fibers, each with a 62.5 $\mu$m core diameter, forming a 375 $\mu$m diameter honeycomb pattern. Each fiber then propagates the collected intensity signal to an independent single-photon avalanche diode, recording the number of detection events. After normalizing the measured counts, this setup associates the sampled polarization state with the corresponding relative frequencies of each detector, which we term a count distribution. For more experimental details, see Sec. 4.1.
@@ 2.1 Experimental and data processing setup
We used the dataset to train a fully-connected deep neural network for reconstructing the polarization state given the associated count distribution, see Sec. 4.3 for details. The optimized and fully-trained network was evaluated on a test set of previously unobserved polarization states, also acquired from the experimental setup. We achieved the unprecedented average infidelity of $8\times 10^{-4}$ with a $[5\times 10^{-5},2\times 10^{-3}]$ confidence interval using the 10th and 90th percentile. These results reflect the highest achievable measurement accuracy as they approach the fidelity of our polarization state preparation. Furthermore, we conducted additional validation of our results using a 1951 USAF test, as detailed in Sec. 4.4. This supplemental analysis provides further insight into the overall performance of our approach.
2.2 Performance evaluation
@@
Furthermore, let us discuss the long-term stability of our approach and its resistance to environmental effects. Despite using a 5 cm long few-mode fiber in an open environment without elaborated fixation, the polarization sensor shows sufficient stability for a week without additional calibration or network retraining. For in-situ and microendoscopic applications, the fiber is typically inserted in a cannula, preventing bending and micro deformations. The stability can be further enhanced using a few-mm or sub-mm long fiber encased within a ceramic or glass ferrule tip. The millimeter-long fiber provides sufficient intermodal coupling, and the encasing fixes the fiber and its position relative to the fiber array. With these modifications, we expect an improvement in stability by up to two orders of magnitude. Employing very short few-mode fiber represents the major difference compared to imaging applications, where long and massively multimode fibers have to be used with deteriorating impact on the stability [22, 23, 24, 25, 26, 27, 9]. In summary, our polarization sensor represents an outstanding instrument for efficient and accurate real-time and high-speed polarimetry applications.
2.3 Applications
@@
Figure 3: Visualization of dense connective tissue: (left) an intensity image, (middle) a scan using the all-fiber polarization sensor, (right) an image using a stand-alone polarization microscope. The resulting Bloch parameters of the all-fiber scan are represented as an RGB false-colored image. The three highlighted pixels characterize the purple and green segments in the polarization structure alongside a reference background polarization. Their respective Bloch parameters are $\text{A}=(0.34,-0.43,0.82),$ $\text{B}=(-0.90,0.05,0.06),$ and $\text{C}=(-0.04,-0.18,0.98).$
We first demonstrate the sensing ability of our approach by conducting a polarization-sensitive scan of dense connective tissue (AmScope PS25W). While the primary application of our sensor significantly differs from polarization imaging, this optically anisotropic specimen is famously known in the polarization microscopy field. Visualizing its typical structures further underscores the accuracy of the presented polarimetric sensor. Using the experimental arrangement outlined in Sec. 4.1, we replaced the calibration polarization preparation stage with the tissue specimen. The few-mode fiber positioned proximal to the tissue scanned over its area, capturing the induced polarization transformation. The reconstructed polarization states were expressed as three Bloch parameters, i.e., renormalized Stokes parameters [44], and visualized as a false-colored RGB image, see Sec. 4.4 for details. In Fig. 3, we present a side-by-side comparison of this polarization scan (middle) with an intensity image (left) captured from an identical tissue region. This intensity image was obtained using a monochrome camera (ImagingSource DMK 23U274) with a pixel size of $4.4\times 4.4~{}\mu$m. We illuminated the specimen with 0.8 $\mu$m light and projected the resulting image onto the camera through a 25.4 mm focal-length lens, providing an estimated fivefold magnification. Moreover, the third panel (right) depicts a polarization image of the same specimen obtained using a stand-alone polarization microscope. Its experimental setup comprises a pair of crossed polarizers and a monochrome camera (ImagingSource DMK 23U274). As illustrated in Fig. 3, both polarization-based methods reveal underlying polarization structures that remain concealed in the intensity profile. This typical inner structure of the dense connective tissue verifies that our sensor performs polarization characterization, providing complete information about the polarization state.
@@
As the dense connective tissue is a static specimen, the complete polarization analysis provided by the all-fiber sensor can also be performed using an advanced imaging polarimeter comprising rotating wave plates. However, such a technique is insufficient for characterizing a rapidly changing polarization state induced by moving or evolving specimens. These scenarios require significantly faster approaches that can be performed in a single-shot regime. We demonstrate the feasibility of our sensor for such applications by conducting a polarization measurement to characterize the floating diatom. These microscopic unicellular organisms comprise a complex inner structure that can exhibit anisotropic properties. By continuously measuring the polarization state, we characterized the polarization transition caused by a moving diatom within the same experimental arrangement as in the previous application. The evolution of the Bloch parameters corresponding to this dynamic process is depicted in Fig. 4 with red lines. Compared to the blue lines of background polarization, the diatom properties are visible in the modulation of Bloch parameters. One could count the number of living specimens, estimate their size and shape, or even classify the detected organisms based on their birefringent properties. The sensor can also find applications for fast in-situ material inspection, local strain analysis, and micro crystal growth monitoring. These showcases underscore the versatility of our polarization sensor for various sensing applications and its valuable contributions to low-light high-speed polarimetry.
@@ 3 Conclusion
We have developed a single-shot polarization sensor based on light propagation through a short rigid piece of a few-mode fiber, resulting in a polarization-dependent scattering. This process provides a coarse speckle pattern with low signal losses and unprecedented long-term stability, which is virtually unaffected by fiber bending. The defining aspect of our approach is the sparse sampling of this pattern using a fiber array instead of capturing the whole intensity profile. This characteristic allows the use of any photonic detectors, including high-speed or single-photon detectors, enabling fast response and operation at low light intensities down to the picowatt level. The implementation of deep learning algorithms in the system ensures an accurate reconstruction, providing complete information about the polarization state, including partial polarization, with an infidelity of $8\times 10^{-4}$. Complemented by the high operation speed, it allows for both real-time and high-repetition-rate polarization sensing with remarkable precision. The all-fiber sensor, featuring a compact design with no moving components, is particularly well-suited for reducing the invasiveness of biomedical procedures in constrained spaces and under low-light conditions. Furthermore, we can enhance the spatial resolution of our sensor by tapering the tip of the few-mode fiber. This method could even provide comparable resolution as near-field scanning optical microscopy while preserving the single-photon polarization sensitivity.
In the view of optically-assisted machine learning, the random propagation in multimode fiber followed by a computational neural network represents an instance of an extreme learning machine or reservoir computing [34, 47]. By demonstrating the single-photon sensing capability of these architectures, we have provided the foundation for decreasing energy consumption in optical sensing, similar to what has been studied recently for computing [48]. Furthermore, to fully quantify the flow of information about the measured quantity (polarization state of light in our case), one may contemplate utilizing the Fisher information concept that has been only recently adapted for wave scattering [49]. The crucially missing part, and possible future focus of the research community, might be the analysis of the flow of Fisher information in neural networks [50, 51]. This would complement the analysis of the physical part of the sensor and allow for the full theoretical information analysis of the deep-learning aided metrology based on random wave scattering.
In summary, our polarization sensor significantly advances the field, answering the demand for a combination of compactness, accuracy, low light sensitivity, and high operational speed, which opens up new possibilities for highly accurate polarization sensing across various applications. Furthermore, our contribution lays down novel research pathways to single-photon level sensing. It also poses new questions on information and confidence evaluation of systems consisting of random media together with deep neural networks. The ability to extract meaningful data from minimal input has far-reaching implications in fields where resource efficiency and sensitivity are paramount.
4 Materials and methods
@@ 4.1 Experimental setup
In addition to the information provided in the main text, we include a detailed scheme of our all-fiber polarization sensor, illustrated in Fig. 5. This scheme comprises three configurations. In the first configuration, the twisted nematic liquid crystal device [45] induces polarization transformation on the attenuated coherent signal generated by a fiber-coupled continuous 810 nm semiconductor laser diode (QPhotonics QFLD). We independently verified the accurate operation of the liquid crystal device in the second configuration by replacing it with a reference polarimeter based on rotating wave plates. The third configuration represents the sensor application and comprises the birefringent specimen under investigation. In all cases, the modified signal is collected by the 5 cm long SMF28 few-mode fiber. The number of modes supported by this fiber is characterized in Sec. 4.2.
@@
Figure 5: A detailed visualization of the experimental setup for the all-fiber sensor. The twisted nematic liquid crystal device, labeled as (a), is present only during dataset acquisition for deep neural network training. In the verification stage, depicted in (b), a reference polarimeter based on wave plates is employed to validate the accurate operation of the liquid crystal device. Finally, this segment is replaced with a polarization-modulating specimen (c) during the USAF test, dense connective tissue, and diatom measurements. The green panel depicts a polarization-dependent speckle pattern generated by the few-mode fiber. The red circles represent the sparse areas sampled by the subsequent fiber array.
@@ 4.2 Supported fiber modes analysis
$V=\frac{2\pi}{\lambda}r_{\text{core}}\text{NA},$ (1)
where $\lambda=810~{}\text{nm}$ is the wavelength of propagated light, $r_{\text{core}}=4.1~{}\mu\text{m}$ is the fiber core radius, and $\text{NA}=0.14$ is the numerical aperture. The calculated value of $V=4.45>2.405$ indicates that the fiber supports more than one mode per polarization direction. Subsequently, the number of modes $M$ is approximately given by the relation
$M\approx\frac{V^{2}}{2}=9.9,$ (2)